A mass m at the end of a spring oscillates with a frequency of 0.92 Hz . When an additional 770 g mass is added to m, the frequency is 0.58 Hz Answer in two significant figures with correct units
Answer #1

as we know

( 2 pi f) ^2 = k/m.. (i)

from above expression

f^2 is inversely proportional to mass

(0.92/0.58)^2 = (770 + m) /m

m = 507.9 g

m = 0.5079 kg

=========

do comment in case any doubt, will reply for sure.. Goodluck